# Chain of Strategy Optimization Makes Large Language Models Better Emotional Supporter

### v1

@@ Abstract
The growing emotional stress in modern society has increased the demand for Emotional Support Conversations (ESC). While Large Language Models (LLMs) show promise for ESC, they face two key challenges: (1) low strategy selection accuracy, and (2) preference bias, limiting their adaptability to users’ emotional needs. Existing supervised fine-tuning (SFT) struggles to address these issues, as it rigidly trains models on single gold-standard responses without modeling nuanced strategy trade-offs. To overcome these limitations, we propose Chain-of-Strategy Optimization (CSO), a novel approach that optimizes strategy selection preferences at each dialogue turn. We first leverage Monte Carlo Tree Search to construct ESC-Pro, a high-quality preference dataset with turn-level strategy-response pairs. Training on ESC-Pro with CSO improves both strategy accuracy and bias mitigation, enabling LLMs to generate more empathetic and contextually appropriate responses. Experiments on LLaMA-3.1-8B, Gemma-2-9B, and Qwen2.5-7B demonstrate that CSO outperforms standard SFT, highlighting the efficacy of fine-grained, turn-level preference modeling in ESC. [^1] Our code and data can be found in supplementary files.
@@
Weixiang Zhao1thanks:    Equal contribution, Xingyu Sui1footnotemark: , Xinyang Han1, Yang Deng2, Yulin Hu1, Jiahe Guo1 Libo Qin3, Qianyun Du4, Shijin Wang4, Yanyan Zhao1, Bing Qin1, Ting Liu1 1Harbin Institute of Technology, 2Singapore Management University 3Central South University, 4iFLYTEK AI Research (Central China), iFLYTEK Co., Ltd {wxzhao, xysui, yyzhao}@ir.hit.edu.cn
@@ 1 Introduction
However, achieving high-quality ESC with LLMs remains a significant challenge due to two core limitations: 1) LLMs struggle with low strategy selection accuracy, often failing to choose contextually appropriate support strategies (Zhao et al., 2023a; Chen et al., 2023; Farhat, 2024). 2) LLMs exhibit strategy preference bias, favoring certain strategies rigidly rather than adapting to users’ dynamic emotional needs (Kang et al., 2024a). Due to the strategy-driven nature of ESC, nuanced trade-offs between strategies are critical (Liu et al., 2021; Deng et al., 2023; Zhao et al., 2023b). While supervised fine-tuning (SFT) is the mainstream training approach (Ouyang et al., 2022; Zheng et al., 2024), its inherent rigidity exacerbates these challenges. By training exclusively on single golden strategies and responses, SFT teaches LLMs which strategies to apply but fails to clarify why certain strategies are inappropriate in specific contexts. This leaves models unable to grasp the contextual nuances required for dynamic adaptation, resulting in less empathetic and supportive responses.
To this end, we propose Chain-of-Strategy Optimization (CSO), a novel approach designed to enhance both strategy selection accuracy and adaptive strategy preference in ESC. CSO builds upon preference learning techniques, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), and we employ Monte Carlo Tree Search (MCTS) to enrich existing ESC datasets (Liu et al., 2021; Zheng et al., 2024) by generating turn-level (strategy, response) preference pairs.
As illustrated in the middle of Figure 1, we first expand gold-standard ESC conversations into a conversation tree, where each layer of nodes represents a round of (strategy, response) pairs, and sibling nodes within the same layer correspond to different strategic responses. During the MCTS process, we design a comprehensive value function to assess newly expanded child nodes in terms of empathy, information, humanoid, and strategy. This ensures that the assigned scores accurately reflect the soundness of the strategies represented in each node. Then, we extract a refined ESC dataset for Preference optimization, ESC-Pro, as shown in the right of Figure 1. Specifically, low-scoring nodes are categorized as non-prefer samples, identifying suboptimal strategies, while both the original gold-standard nodes and newly discovered high-scoring nodes are included as prefer samples, enhancing the dataset with high-quality strategy examples. Training on ESC-Pro with CSO not only steers the model toward selecting superior strategies over weaker alternatives at each dialogue turn but also expands the diversity of possible conversation paths, mitigating preference biases.
@@ 2 Related Work
Emotional Support Conversations (ESC) Liu et al. (2021) center around interactions between a user, referred to as the seeker, who is experiencing emotional distress, and a supporter, whose goal is to alleviate the seeker’s emotional intensity by utilizing a set of strategies to guide the conversation. Various approaches have been proposed to build the ESC systems, such as global-to-local hierarchical graph network Peng et al. (2022), incorporating commonsense knowledge Tu et al. (2022), and modeling emotions and sematics Zhao et al. (2023b). With the development of LLMs, some works aim to directly leverage the performance of these models without altering their architecture. Liu et al. (2023) apply SFT to the LLaMA-7B model for the ESC task and introduce ChatCounselor, a model specialized in ESC tasks that outperforms general-purpose models. Chen et al. (2023) and Qiu et al. (2023) expand single-turn empathic responses into multi-turn dialogues and performed fine-tuning on a high-quality ESC dataset they constructed, thereby improving the model’s ESC performance.
Figure 2: The overall framework of CSO. (a) Dialogue Expansion with MCTS: We leverage Monte Carlo Tree Search (MCTS) to systematically explore and refine ESC strategies by constructing a structured conversation tree. (b) Preference Data Construction: We extract high-quality strategy preference pairs from the expanded conversation tree to form the ESC-Pro dataset. (c) Chain-of-Strategy Optimization (CSO): We fine-tune LLMs on ESC-Pro using preference optimization techniques to enhance strategy selection accuracy and mitigate preference bias.
@@ 3 Methodology
We introduce Chain-of-Strategy Optimization (CSO), a novel approach that refines strategy selection at each dialogue turn through structured preference modeling. As shown in Figure 2 This optimization paradigm consists of three key stages: (1) Dialogue Expansion with MCTS, (2) Preference Data Construction and (3) Chain-of-Strategy Optimization (CSO).
@@ 3.1 Dialogue Expansion with MCTS
To enrich exists ESC dataset with turn-level preference annotation, we employ Monte Carlo Tree Search (MCTS) to construct a structured conversation tree, where each node represents a specific dialogue state. The search process iteratively refines dialogue strategies through four key stages: selection, expansion, simulation, and backpropagation. This enables the model to explore various strategy paths and identify optimal responses based on predefined value function.
@@ 3.1.4 Simulation
where $E$, $I$, $H$, and $S$ are calculated based on the conversation history at each step. $\alpha$ is a scaling hyper-parameter. A bias $b$ is introduced to adjust the reward, allows the system to treat rewards lower than $-b$ as negative, helping guide the search towards higher-scoring nodes. The reward for the simulation is averaged over all Supporter turns during the simulation and is used to update the node’s Q-value in the backpropagation phase:
@@ 3.2 ESC-Pro Preference Data Construction
[…]
@@
$\mathcal{D}=\bigcup_{P\in\mathcal{P}}\left\{\begin{aligned} &\{(S_{w},S_{l})\mid S_{w}\in P,S_{l}\in P,\\ &\quad Q(S_{w})>\theta,Q(S_{l})<\theta\}\end{aligned}\right.$ (6)
where $\mathcal{P}$ denotes the set of all valid paths. The dataset $\mathcal{D}$ contains all preference pairs $(S_{w},S_{l})$ extracted from valid paths. By incorporating both preferred and non-preferred strategies, ESC-Pro provides a rich training signal, allowing LLMs to learn nuanced strategy trade-offs and improve adaptive decision-making.
@@ 4.3 Strategy Analysis
The ExTES dataset originally contains 16 distinct support strategies. To reduce the search space and improve computational efficiency, we merge similar strategies into 8 categories (see Appendix D for details). We analyze the distribution of these strategies across six conversation stages. Given a dialogue with $N$ utterances, each utterance $k$ (where $k=1,…,N$) is assigned to stage $i$ using:
@@ 5.1 Experimental Setup
We evaluate ESC performance using the ExTES test set and a strategy test set constructed following Kang et al. (2024b). The evaluation focuses on two four metrics: strategy selection accuracy, measured by macro F1 score $\mathcal{Q}$ and weighted F1 score $\mathcal{Q_{W}}$, and strategy preference bias $\mathcal{B}$, which quantifies the deviation between the model’s selected strategies and the ideal strategy preferences. And ROUGE-L (R-L) (Lin, 2004) is also adopted to evaluate the semantic aspects of the generated response.
@@
Given that most recent methods to improving ESC capabilities in LLMs rely on supervised fine-tuning (SFT) (Chen et al., 2023; Qiu et al., 2023; Zheng et al., 2024), we compare CSO against LoRA-based (Hu et al., 2022) and full-parameter SFT models trained on ESC-Pro-SFT. Additionally, we evaluate several decoding strategies, including Direct-Refine, Self-Refine, and the inclusion of examples in the prompt, to assess their impact on model performance. Please refer to Appendix E for detailed description.
@@ 5.2.1 Automatic Evaluation Results
Table 2 demonstrates the performance comparison of CSO and baselines based on LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct and gemma-2-9B-it.
CSO improves strategy accuracy while reducing bias. Table 2 shows that CSO significantly enhances strategy accuracy while effectively reducing strategy bias. While SFT without preference optimization slightly improves strategy selection accuracy, it also increases strategy bias, limiting adaptability. In contrast, CSO mitigates bias while boosting accuracy, highlighting the necessity of preference optimization in ESC strategy selection. Our results also highlight the limitations of decoding-based approaches, which fail to achieve stable improvements, often leading to lower strategy accuracy and increased bias.
CSO enhances ESC performance across different models. CSO improves ESC performance across all backbone models, regardless of their initial capabilities. It enhances strategy accuracy and reduces bias in both weaker models (Qwen2.5-7B-Instruct) and stronger models (Gemma-2-9B-it), demonstrating its versatility and robustness.
@@ 5.2.2 Human Evaluation Results
Table 3: Human evaluation results comparing CSO and SFT. Win indicates cases where CSO-generated responses were preferred, while Lose represents cases where SFT responses were rated higher. Tie indicates no preference.
The results in Table 3 show that CSO consistently outperforms SFT across all human evaluation metrics. Specifically: CSO achieves higher Acceptance (68.00% win rate), Effectiveness (58.33%), and Sensitivity (60.67%), indicating that its responses are more appropriate, impactful, and emotionally attuned. Inter-rater agreement ($\kappa$ scores between 0.55–0.65) indicates a moderate to high level of consistency among human evaluators. These findings further validate that preference optimization enhances ESC performance, making responses more empathetic and aligned with user needs.
@@ 5.3 Ablation Study
We fine-tune the model using LoRA with the same hyperparameters as ESC-Pro and compare ESC-Pro-Random with both ESC-Pro and standard SFT. As shown in Table 4, ESC-Pro-Random performs slightly better than SFT but remains consistently inferior to ESC-Pro across all bakbone models. This confirms that our method of leveraging low-scoring nodes searched and verified by MCTS as non-preferred data is both effective and meaningful, reinforcing the importance of structured preference learning in ESC.
@@ 5.4 Data Volume Analysis
We analyze the impact of data volume on model performance by varying the number of ESC-Pro dialogues used for fine-tuning. We experiment with 0 (baseline), 100, 200, 300, 400, and the full set (423 dialogues), applying LoRA fine-tuning under consistent experimental settings. Figure 4 presents the results, with SFT performance indicated by dashed lines for reference.
@@ 5.5 Analysis of Different Preference Optimization Algorithms
To further validate the efficacy of CSO, we examine the impact of integrating alternative preference learning methods into our framework. We replace the default DPO with various existing alternatives, including IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024) and ORPO (Hong et al., 2024), and evaluate their performance. The results consistently show that all preference optimization variants outperform standard SFT, reinforcing the advantages of preference-driven learning in ESC. For detailed results and discussion, please refer to Appendix G.
@@ 6 Conclusion
In this work, we address key challenges in LLM-driven ESC, particularly low strategy selection accuracy and preference bias. To tackle these issues, we propose Chain-of-Strategy Optimization (CSO), a novel approach that optimizes strategy selection preferences at each dialogue turn. CSO systematically expands existing ESC datasets with Monte Carlo Tree Search, constructing a conversation tree where different strategy-response pairs are evaluated to generate ESC-Pro, a refined preference dataset. By training LLMs on ESC-Pro, CSO improves strategy accuracy, reduces bias toward fixed strategies, and enhances adaptability to user emotions. Extensive experiments on LLaMA-3.1-8B-Instruct, Gemma-2-9B-it, and Qwen2.5-7B-Instruct demonstrate that CSO significantly outperforms standard SFT and decoding-based methods. These results validate the effectiveness of turn-level preference modeling in improving ESC quality.
@@
Despite the effectiveness of CSO in improving strategy selection and preference alignment, our study has several limitations: Due to limited computational resources, our experiments were conducted on mid-scale LLMs (7B–9B parameters). While these models are representative, larger-scale models (e.g., 30B+) could further enhance performance and provide deeper insights into CSO’s scalability. Our ESC-Pro dataset was generated using a seed set of 100 dialogues, expanded through MCTS-based search. While the results demonstrate significant performance gains, a larger seed dataset or alternative expansion strategies (e.g., human-in-the-loop validation) could further enhance the diversity and quality of preference data.

### v3

@@ Abstract
The growing emotional stress in modern society has increased the demand for Emotional Support Conversations (ESC). While Large Language Models (LLMs) show promise for ESC, they face two key challenges: (1) low strategy selection accuracy, and (2) preference bias, limiting their adaptability to users’ emotional needs. Existing supervised fine-tuning (SFT) struggles to address these issues, as it rigidly trains models on single gold-standard responses without modeling nuanced strategy trade-offs. To overcome these limitations, we propose a novel two-stage framework that optimizes strategy selection preferences at each dialogue turn. We first leverage Monte Carlo Tree Search to construct ESC-Pro, a high-quality preference dataset with turn-level strategy-response pairs. Then training on ESC-Pro with Chain-of-Strategy Optimization (CSO) improves both strategy accuracy and bias mitigation, enabling LLMs to generate more empathetic and contextually appropriate responses. Experiments on LLaMA-3.1-8B, Gemma-2-9B, and Qwen2.5-7B demonstrate that CSO outperforms standard SFT, highlighting the efficacy of fine-grained, turn-level preference modeling in ESC. [^1] Our code and data can be found in https://github.com/XingYuSSS/CSO.
@@
Weixiang Zhao1thanks: Equal contribution, Xingyu Sui1footnotemark: , Xinyang Han1, Yang Deng2, Yulin Hu1, Jiahe Guo1 Libo Qin3, Qianyun Du4, Shijin Wang4, Yanyan Zhao1thanks: Corresponding author, Bing Qin1, Ting Liu1 1Harbin Institute of Technology, 2Singapore Management University 3Central South University, 4iFLYTEK AI Research (Central China), iFLYTEK Co., Ltd {wxzhao, xysui, yyzhao}@ir.hit.edu.cn
@@ 1 Introduction
However, achieving high-quality ESC with LLMs remains a significant challenge due to two core limitations: 1) LLMs struggle with low strategy selection accuracy, often failing to choose contextually appropriate support strategies (Zhao et al., 2023b; Chen et al., 2023; Farhat, 2024). 2) LLMs exhibit strategy preference bias, favoring certain strategies rigidly rather than adapting to users’ dynamic emotional needs (Kang et al., 2024). Due to the strategy-driven nature of ESC, nuanced trade-offs between strategies are critical (Liu et al., 2021; Deng et al., 2023; Zhao et al., 2023c). While supervised fine-tuning (SFT) is the mainstream training approach (Ouyang et al., 2022; Zheng et al., 2024), its inherent rigidity exacerbates these challenges. By training exclusively on single golden strategies and responses, SFT teaches LLMs which strategies to apply but fails to clarify why certain strategies are inappropriate in specific contexts. This leaves models unable to grasp the contextual nuances required for dynamic adaptation.
To tackle these challenges, we advocate for a more fine-grained and turn-level approach to optimizing ESC. Specifically, we propose a two-stage framework that first constructs rich preference data at the (strategy, response) level, and then refines model behavior through preference learning.
In the first stage, we utilize Monte Carlo Tree Search (MCTS) to expand gold-standard ESC conversations into a conversation tree, where each layer represents a dialogue turn, and sibling nodes correspond to different strategic responses. As illustrated in the middle of Figure 1, we design a comprehensive value function that evaluates each response along four key dimensions—Empathy, Information, Humanoid, and Strategy—to ensure high-quality exploration. From this tree, we extract a refined ESC dataset for Preference optimization, ESC-Pro, as shown in the right of Figure 1. Specifically, low-scoring nodes are categorized as non-prefer samples, identifying suboptimal strategies, while both the original gold-standard nodes and newly discovered high-scoring nodes are included as prefer samples, enhancing the dataset with high-quality strategy examples.
In the second stage, we introduce Chain-of-Strategy Optimization (CSO), a preference optimization approach that explicitly targets strategy-level improvements. As shown in the right of Figure 1, training on ESC-Pro with CSO enables LLMs to not only select better strategies over weaker ones at each dialogue turn, but also to explore diverse conversation paths and avoid rigid preference patterns. This chain-like optimization across turns improves both local adaptability and long-range strategic coherence.
@@ 2 Related Work
Emotional Support Conversations (ESC) Liu et al. (2021) center around interactions between a user, referred to as the seeker, who is experiencing emotional distress, and a supporter, whose goal is to alleviate the seeker’s emotional intensity by utilizing a set of strategies to guide the conversation. Various approaches have been proposed to build the ESC systems, such as global-to-local hierarchical graph network Peng et al. (2022), incorporating commonsense knowledge Tu et al. (2022), and modeling emotions and semantics Zhao et al. (2023a, c). With the development of LLMs, some works aim to directly leverage the performance of these models without altering their architecture. Liu et al. (2023) apply SFT to the LLaMA-7B model for the ESC task and introduce ChatCounselor, a model specialized in ESC tasks that outperforms general-purpose models. Chen et al. (2023) and Qiu et al. (2023) expand single-turn empathic responses into multi-turn dialogues and performed fine-tuning on a high-quality ESC dataset they constructed, thereby improving the model’s ESC performance.
Figure 2: The overall framework. (a) Dialogue Expansion with MCTS: We leverage Monte Carlo Tree Search (MCTS) to systematically explore and refine ESC strategies by constructing a structured conversation tree. (b) Preference Data Construction: We extract high-quality strategy preference pairs from the expanded conversation tree to form the ESC-Pro dataset. (c) Chain-of-Strategy Optimization (CSO): We fine-tune LLMs on ESC-Pro using preference optimization techniques to enhance strategy selection accuracy and mitigate preference bias.
@@ 3 Methodology
We introduce a novel framework that refines strategy selection at each dialogue turn through structured preference modeling. As shown in Figure 2, this optimization paradigm consists of three key components: (1) Dialogue Expansion with MCTS, (2) Preference Data Construction and (3) Chain-of-Strategy Optimization (CSO).
@@ 3.1 Dialogue Expansion with MCTS
To enrich exists ESC dataset with turn-level preference annotation, we employ Monte Carlo Tree Search (MCTS) to construct a structured conversation tree, where each node represents a specific dialogue state. The search process iteratively refines dialogue strategies through four key stages: selection, expansion, simulation, and backpropagation. This enables the model to explore various strategy paths and identify optimal responses based on a predefined value function.
@@ 3.1.4 Simulation
where $E$, II, $H$, and $S$ are calculated based on the conversation history at each step. $\alpha$ is a scaling hyper-parameter. A bias $b$ is introduced to adjust the reward, allows the system to treat rewards lower than b-b as negative, helping guide the search towards higher-scoring nodes. The reward for the simulation is averaged over all Supporter turns during the simulation and is used to update the node’s Q-value in the backpropagation phase:
@@ 3.2 ESC-Pro Preference Data Construction
$\text{is\_end\_node}(S_{L})\ \;\text{and}\ \;Q(S_{i})>\theta\ \;\forall S_{i}\in P$ (5)
Here, $\text{is\_end\_node}(S_{L})$ ensures that the last node in the path, $S_{L}$, is a termination node, and $Q(S_{i})>\theta$ ensures that all nodes in the path meet the quality threshold.
@@
$\mathcal{D}=\bigcup_{P\in\mathcal{P}}\left\{\begin{aligned} &\{(S_{w},S_{l})\mid S_{w}\in P,S_{l}\in\text{Siblings}(S_{w}),\\ &\quad Q(S_{w})>\theta,Q(S_{l})<\theta\}\end{aligned}\right$. (6)
where $\mathcal{P}$ denotes the set of all valid paths, $\text{siblings}(S_{w})$ denotes all sibling nodes of the node $S_{w}$. The dataset $\mathcal{D}$ contains all preference pairs $(S_{w},S_{l})$ extracted from valid paths. By incorporating both preferred and non-preferred strategies, ESC-Pro provides a rich training signal, allowing LLMs to learn nuanced strategy trade-offs and improve adaptive decision-making.
@@ 4.3 Strategy Analysis
The ExTES dataset originally contains 16 distinct support strategies. To reduce the search space and improve computational efficiency, we merge similar strategies into 8 categories (see Appendix D for details). We analyze the distribution of these strategies across six conversation stages. Given a dialogue with $N$ utterances, each utterance $k$ (where k=1,,Nk=1,…,N) is assigned to stage $i$ using:
@@ 5.1 Experimental Setup
We evaluate ESC performance on the ExTES test set and a strategy test set constructed following Kang et al. (2024). The evaluation includes four metrics: strategy selection accuracy (macro F1 $\mathcal{Q}$, weighted F1 $\mathcal{Q_{W}}$), strategy preference bias ($\mathcal{B}$), which measures deviation from ideal strategy distributions, and ROUGE-L (R-L) (Lin, 2004) for assessing response semantics.
@@
We compare CSO with both LoRA-based (Hu et al., 2022) and full-parameter supervised fine-tuning (SFT) models trained on ESC-Pro-SFT, following recent ESC approaches (Chen et al., 2023; Qiu et al., 2023; Zheng et al., 2024). We also evaluate decoding strategies such as Direct-Refine, Self-Refine, and in-context example prompting. In addition, we benchmark against strong proprietary models, including GPT-4o, Gemini-2.0, and the Claude series. See Appendix E for details.
@@ 5.2 Overall Results
GPT-4o-mini 35.68 2.65 42.08 24.91
GPT-4o 27.98 2.65 39.55 24.26
Gemini-2.0-Pro 27.00 1.90 46.59 22.71
Claude-3.5-Sonnet 20.97 1.55 41.00 19.82
Claude-3.7-Sonnet 31.50 1.18 48.13 21.01
@@ 5.2.1 Automatic Evaluation Results
Table 2 demonstrates the performance of CSO and baselines based on LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma-2-9B-it.
CSO improves strategy accuracy while reducing bias. Table 2 shows that CSO significantly enhances strategy accuracy while effectively reducing strategy bias. While SFT without preference optimization slightly improves strategy selection accuracy, it also increases strategy bias, limiting adaptability. In contrast, CSO mitigates bias while boosting accuracy, highlighting the necessity of preference optimization in ESC strategy selection. Our results also highlight the limitations of decoding-based approaches, which fail to achieve stable improvements, often leading to lower strategy accuracy and increased bias. Notably, CSO-equipped open-weight models outperform several leading closed-source models on both accuracy and bias, demonstrating its strong generalization.
CSO enhances ESC performance across different models. CSO improves ESC performance across all backbone models. It enhances strategy accuracy and reduces bias in both weaker models (Qwen2.5-7B-Instruct) and stronger models (Gemma-2-9B-it), demonstrating its versatility and robustness. Moreover, we further evaluate CSO on a larger model, Qwen2.5-32B, and observe consistent improvements, confirming the scalability of our approach. Detailed results and analysis are provided in Appendix G.2.
@@ 5.2.2 Human Evaluation Results
Table 3: Human evaluation comparing CSO and SFT. Win indicates CSO-generated responses are preferred, while Lose represents cases where SFT responses are rated higher. Tie indicates no preference.
Results in Table 3 show that CSO consistently outperforms SFT across all human evaluation metrics. Specifically, CSO achieves higher Acceptance (68.00% win rate), Effectiveness (58.33%), and Sensitivity (60.67%), indicating that its responses are more appropriate, impactful, and emotionally attuned. Inter-rater agreement ($\kappa$ scores between 0.55–0.65) indicates a moderate to high level of consistency among evaluators. These findings further validate that preference optimization enhances ESC performance, making responses more empathetic and aligned with user needs. More details of human evaluation can be found in I.
@@ 5.3 Ablation Study
We fine-tune the model using LoRA with the same hyperparameters as ESC-Pro and compare ESC-Pro-Random with both ESC-Pro and standard SFT. As shown in Table 4, ESC-Pro-Random performs slightly better than SFT but remains consistently inferior to ESC-Pro across all backbone models. This confirms that our method of leveraging low-scoring nodes searched and verified by MCTS as non-preferred data is both effective and meaningful, reinforcing the importance of structured preference learning in ESC.
@@ 5.4 Data Volume Analysis
We analyze the impact of data volume by varying the number of ESC-Pro dialogues used for fine-tuning. We experiment with 0, 100, 200, 300, 400, and the full set (423 dialogues), applying LoRA fine-tuning under consistent experimental settings. Figure 4 presents the results, with SFT performance indicated by dashed lines for reference.
@@ 5.5 Analysis of Different Preference Optimization Algorithms
To further validate the efficacy of CSO, we examine the impact of integrating alternative preference learning methods into our framework. We replace the default DPO with various existing alternatives, including IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024) and ORPO (Hong et al., 2024), and evaluate their performance. The results consistently show that all preference optimization variants outperform standard SFT, reinforcing the advantages of preference-driven learning in ESC. For detailed results and discussion, please refer to Appendix G.1.
@@ 6 Conclusion
In this work, we address low strategy selection accuracy and preference bias challenges in ESC. We propose a novel two-stage framework that optimizes strategy selection preferences at each dialogue turn. We first expand existing ESC datasets with Monte Carlo Tree Search, constructing a conversation tree where different strategy-response pairs are evaluated to generate ESC-Pro, a refined preference dataset. By training LLMs on ESC-Pro, Chain-of-Strategy Optimization (CSO) improves strategy accuracy, reduces bias, and enhances adaptability to user emotions. Extensive experiments on LLaMA-3.1-8B-Instruct, Gemma-2-9B-it, and Qwen2.5-7B-Instruct demonstrate that CSO significantly outperforms standard SFT and decoding-based methods, validating the efficacy of turn-level preference modeling in improving ESC quality.
@@
Despite the effectiveness of CSO in improving strategy selection and preference alignment, our study has several limitations: Due to limited computational resources, our experiments were conducted on mid-scale LLMs (7B–32B parameters). While these models are representative, larger-scale models (e.g., 70B+) could further enhance performance and provide deeper insights into CSO’s scalability. Our ESC-Pro dataset was generated using a seed set of 100 dialogues, expanded through MCTS-based search. While the results demonstrate significant performance gains, a larger seed dataset or alternative expansion strategies (e.g., human-in-the-loop validation) could further enhance the diversity and quality of preference data. Future work should also place more emphasis on personalization (Liu et al., 2025; Qiu et al., 2025b, a) and safety (Zhao et al., 2025) to ensure that CSO-driven dialogue systems are not only effective but also aligned with user-specific needs and robust against potential risks.
@@
Acknowledgments
We thank the anonymous reviewers for their comments and suggestions. This work was supported by the New Generation Artificial Intelligence-National Science and Technology Major Project 2023ZD0121100, the National Natural Science Foundation of China (NSFC) via grant 62441614 and 62176078, the Fundamental Research Funds for the Central Universities, and the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C004).